(h-50)x+(h+50)y=0


1
Expert's answer
2021-05-07T09:06:05-0400

h-50)x+(h+50)y=0

hx-50x+hy+50y=0

h(x+y)-50(x-y)=0

Suppose x=y, then

h="\frac{o}{x+y}"